U.S. flag

An official website of the United States government

Dot gov

Official websites use .gov
A .gov website belongs to an official government organization in the United States.

HTTPS

Secure .gov websites use HTTPS
A lock () or https:// means you’ve safely connected to the .gov website. Share sensitive information only on official, secure websites.

Delek Pipeline - E14612

All POL/SITREP's for this site Delek Pipeline - E14612
Haynesville, LA - EPA Region VI
POLREP #6
Printer Friendly  |   PDF
 
U.S. ENVIRONMENTAL PROTECTION AGENCY
POLLUTION/SITUATION REPORT
Delek Pipeline - E14612 - Removal Polrep
Final Removal Polrep

EPA Emergency Response

UNITED STATES ENVIRONMENTAL PROTECTION AGENCY
Region VI

Subject: POLREP #6
Delek Pipeline - E14612

Haynesville, LA
Latitude: 32.9703500 Longitude: -92.9367500


To:
From: Mike McAteer, OSC
Date: 7/14/2014
Reporting Period:

1. Introduction
  1.1 Background
   
Site Number: V6PW    Contract Number:  
D.O. Number:      Action Memo Date:  
Response Authority: OPA    Response Type: Emergency
Response Lead: PRP    Incident Category: Removal Action
NPL Status: Non NPL    Operable Unit:
Mobilization Date: 4/25/2014    Start Date: 4/25/2014
Demob Date: 5/15/2014    Completion Date: 6/10/2014
CERCLIS ID:    RCRIS ID:
ERNS No.:    State Notification:
FPN#: E14612    Reimbursable Account #:

1.1.1 Incident Category

1.1.2 Site Description

On April 25, 2014, Delek Logistics made a report to the National Response Center (NRC No. 1080818), of an estimated 400 bbl oil spill near Haynesville, LA. Delek is the owner/operator of the pipeline and the Responsible Party (RP) for the spill.  The spill was the result of a leak in an 8-inch pipeline that transports crude oil from lease operators in North Louisiana to the Lion Oil Refinery in  El Dorado, AR.  The cause of the leak was reported to be corrosion.  The oil spill impacted a low-land area adjacent to the pipeline and migrated east-northeast into an intermittent creek.  The spill scene is located approximately 2.4 miles northwest of the intersection of Louisiana Alternate Highway (LA Alt Hwy) 2 and Parish Road 179 (Tom Sale Road).

 

1.1.2.1 Location
The discharge occurred approximately 0.3 mile southwest of Holly Ridge Road, within the pipeline right-of-way at Latitude 32.97035 North and Longitude 92.93675 West.  The spill scene is bordered by undeveloped woodland to the north, south, east and west.  Kisatche National Forest is located 1.5 miles south of the spill scene.  The nearest resident is located approximately 0.15 mile south of the spill area.

1.1.2.2 Description of Threat

Oil discharged from the pipeline and flowed down gradient into an unnammed intermittent creek.   The oil is currently contained within a 0.5 mile section of the creek.  From the spill origin, the unnamed creek flows eastward approximately 3500 feet to Doe Branch Creek.  Doe Branch Creek flows east northeast approximately 1.3 miles into Little Corney Creek, which flows approximately 8.6 miles into  Corney Lake.      

1.1.3 Preliminary Removal Assessment/Removal Site Inspection Results

2. Current Activities
  2.1 Operations Section
    2.1.1 Narrative

At 1300 hrs on April 25, EPA Region 6 Phone Duty Officer, Chris Ruhl, activated On Scene Coordinator Mike McAteer and EPA's START-3 contractor (Weston Solutions, Inc.) to mobilize to the incident.  EPA arrived at the incident at 1830 on April 25, 2014 and met with RP representatives to document oil spill containment and cleanup activities.  Shortly before arriving on scene, the  U.S Fish and Wildlife representative for Louisiana contacted OSC McAteer to inform him that there were no known endangered or protected species in the spill area.


2.1.2 Response Actions to Date

On April 25, 2014, the RP mobilized B&P Enterprises, Inc. of Texarkana, Texas to conduct oil spill containment and cleanup operations.  Approximately 20 B&P and 14 Delek Logistics personnel were utilized to conduct containment and initiate recovery operations. Three underflow dams were constructed to contain the downstream migration of oil and provide for oil recovery locations.  Containment hard boom was also deployed approximately 100 feet downstream of the last underflow dam.  Oil recovery operations were initiated and approximately 60 bbl of oil/water mixture was recovered. 

As of April 26, 2014, the oil continues to be contained by the underflow dams within the intermittent creek.  Current cleanup operations consist of removing oily vegetation along the spill pathway and using vacuum trucks to collect areas of pooled oil.  The RP is also releasing water into the creek at the spill origin to wash down the oil and facilitate oil recovery activities.  A Louisiana Department of Wildlife and Fisheries (LDWF) representative has also mobilized to the spill scene to investigate oil impacted wildlife.  It has been reported that an oil stained bird (1) has been observed and captured.  In addition the following dead oil impacted wildlife were observed: snakes (2), fish (1), and a frog (1).   Two representatives from the Louisiana Oil Spill Coordinator's Office are also onsite overseeing cleanup operations.  An additional 70 bbls of oil/water were recovered today.

As of April 27, 2014, the oil impacted area continues to be contained by the underflow dams within the intermittent creek. Cleanup operations continue to consist of removing oily vegetation along the spill pathway, oil absorbent pads placed into oil saturated areas, and vacuum trucks to collect areas of pooled oil. The RP also continues to release water at the spill origin as well as at the upstream underflow dam (No. 1) to facilitate oil recovery activities.  During this reporting period, 97 bags of oily debris and 5 bags of oily sorbent pad were collected, bagged and placed in a roll-off box.  In addition, 265 bbls of oil/water mixture were recovered today.  Total recovery to date is reported to be 460 bbls.   

Wildlife Response Services representatives have mobilized to the site to conduct rehabilitation of oil impacted wildlife. LDWF personnel continue to inspect the oil spill area for impacted wildlife.  LDWF reported that the oil stained bird (1) captured yesterday was cleaned and, in the process of being released, it died.  A bronze frog (1) has been observed, captured, cleaned and released. To date, the following oil-impacted wildlife has been observed: Snakes (1 dead, 1 alive), fish (6 dead), birds (1 dead) and frogs (3 alive).

As of April 28, 2014, the oil impacted area continues to be contained by the underflow dams within the intermittent creek. Cleanup operations continue to consist of removing oily vegetation along the spill pathway, oil absorbent pads placed into oil saturated areas, and vacuum trucks to collect areas of pooled oil. The RP also continues to release water at the spill origin as well as at the upstream underflow dam (No. 1) to facilitate oil recovery activities. The RP contractor brought in an additional 25 persons bringing the total to 47 B&P personnel on site. During this reporting period, 345 bags of oily debris and14 bags of oily sorbent pad were collected, bagged and placed in a roll-off box.  In addition, 75 bbls of oil/water mixture were recovered today though today’s mix was mostly water.  Total recovery to date is reported to be 535 bbls.

LDWF personnel continue to inspect the oil spill area for impacted wildlife.  LDWF collected an oil stained box turtle and are in the process of rehabbing and prepping for release. They also collected an oil stained 5 Lined Skink that has been cleaned and released. To date, the following oil-impacted wildlife has been observed: Turtle (1 alive), Skink (1 alive), snakes (2 dead, 1 alive), fish (8 dead), birds (1 dead) and frogs (3 alive). LDWF observed an unknown turtle near underflow dam (No. 1), but have not been able to capture.

As of May 1, 2014, the oil impacted area continues to be contained by the underflow dams within the intermittent creek. Cleanup operations continue to consist of removing oily vegetation along the spill pathway, oil absorbent pads and booms placed into oil saturated areas, and vacuum trucks to collect areas of pooled oil. The RP also continues to release water at the spill origin to facilitate oil recovery activities. The vast majority of free oil still on the creek is in Sector 1B (see site map under "Documents" section of website).  Also, EPA, LOSCO, and the RP have identified approximately 9 areas adjacent to the creek, primarily in Sector 1B, that are heavily oiled and are acting a source, or potential future source, to further contamination in the creek.  The RP has proposed a plan for these pockets of contaminated soil adjacent to the creek that primarily involves isolating the pockets using earthen berms and then excavating the contaminated zones and backfilling with clean soil and re-seeding.  This work is planned to start tomorrow.  The RP is continuing to use two contractors to aid the oil cleanup. There are currently 110 B&P personnel on site and 15 Code 3 Services personnel onsite.  EPA has increased the FPN to a ceiling of $50,000 for this response.

LDWF personnel continue to inspect the oil spill area for impacted wildlife. LDWF collected one live rat snake (unoiled) that has now been relocated. To date, the following oil-impacted wildlife has been observed: Turtle (1 alive), Skink (1 alive), snakes (2 dead, 2 alive), fish (8 dead), birds (1 dead) and frogs (4 alive).

As of June 10, 2014, the EPA team performed a walk through of the site with LOSCO and LDWF. Site conditions at the spill site are dramatically improved. The oil impacted area continues to be contained by the underflow dams within the intermittent creek. Cleanup operations continue to consist of boom maintenance operations along the spill pathway. The RP has demobed 2 contractors and continues to utilize the remaining contractor for boom mainenance operations. There are currently 7 Code 3 personnel on site  2-3 days a week. The RP had been given approval to remove weir 3 and 4 as weather permitted but has not yet been completed. The RP continues to release water at the spill origin to facilitate oil recovery activities. Minimal sheen was observed in patches along the creek. The Areas of Concern (AOCs) have been excavated, backfilled, and seeded and the damaged pipeline has been repaired. At this time no further site visits are planned.

Weather conditions remain favorable. Sunny, with a high near 89°F with an overnight low expected of 74°F. 

2.1.3 Enforcement Activities, Identity of Potentially Responsible Parties (PRPs)

The Responsible Partly associated with the spill scene is Delek Logistics, 1001 School St, El Dorado, Arkansas.  EPA OSC Mike McAteer submitted a Notice of Federal Interest (NOFI) letter to the on-site RP representative on April 26, 2014.


2.1.4 Progress Metrics

Waste Stream Medium Quantity Manifest # Treatment Disposal
           
           
           


  2.2 Planning Section
    2.2.1 Anticipated Activities

2.2.1.1 Planned Response Activities

The RP will continue to conduct spot checks in the creek over the next few months and address any substantial oil areas.  

2.2.1.2 Next Steps


2.2.2 Issues

  2.3 Logistics Section
    No information available at this time.

  2.4 Finance Section
    No information available at this time.

  2.5 Other Command Staff
    No information available at this time.

3. Participating Entities
  3.1 Unified Command

FOSC - McAteer, US EPA

SOSC - Viator, LOSCO

RP - Scott Wiggins, Delek

3.2 Cooperating Agencies

LDEQ

Louisiana Wildlife and Fisheries

USDOI

Contractors

Weston Solutions

Lion Inc.
B&P
CTEH
Code 3

4. Personnel On Site
  No information available at this time.

5. Definition of Terms
  No information available at this time.

6. Additional sources of information
  No information available at this time.

7. Situational Reference Materials
  No information available at this time.